OpenStudy (anonymous):

Solve the equation 3/2 - 1/x-4 = -2/2x-8 . Use complete sentences to explain how you solved this equation.

OpenStudy (cwrw238):

first you find the LCD of 2, x-4 and 2x-8 do you know why this is?

OpenStudy (anonymous):

is it 2 ?

OpenStudy (anonymous):

hello ? lol

OpenStudy (cwrw238):

no LCD is the lowest common denominate r that means the lowest term into which each term will divide here it is 2x - 8 2x - 8 / 2 = x -4 2x - 8 / x - 4 = 2 and of course 2x - 8 / 2x - 8 = 1

OpenStudy (anonymous):

so all the terms will go into 2x-8 ?

OpenStudy (cwrw238):

all the denominators will so we multiply through by 2x - 8:-